# Non-Hermitian entanglement dip from scaling-induced exceptional criticality

### v1

@@ Abstract
which, unlike in conventional critical scenarios of Figs. 1(b,c), diverges due to $L\rightarrow L_{c}$ rather than vanishing momentum spacings. This square-root divergence shows up in the $\bar{P}$ spectrum near $L\approx L_{c}$ [Fig. 3(a), also see I] which in turn translates to a characteristic $S\sim 1/\sqrt{|L_{c}-L|}$ divergence in the EE which we call the entanglement dip [Fig. 3(b)]  [^10] Since the eigenvalues of the projector $P$ are strictly 00 or $1$, entanglement can be regarded as a measure of how much the entanglement cut i.e. $\bar{\mathcal{R}}$ causes $\bar{P}=\bar{\mathcal{R}}P\bar{\mathcal{R}}$ to depart from being a true projector onto the occupied bands (which extend across the cut), and spectacularly so for negative entanglement.
While such entanglement dips can in principle be infinitely deep, in practice its depth is limited by the how closely $L_{c}$ approaches an integer. Shown in Fig. 3(c) is the minimal entanglement $S_{\text{min}}$ as the hopping asymmetry $t_{L}/t_{R}$ and inter-chain coupling $\delta$ are varied. For each lattice size $L$, the EE dip becomes particularly deep along curves in parameter space (dark pink), even reaching $S_{\text{min}}\approx-3$ in the region shown.
Figure 4: Entanglement dips and scale-dependent GBZ in more general models satisfying Eq. 7. Shown are the $L$-dependent GBZs for illustrative (a) $H=(t_{L}+1/z)\sigma_{+}+(t_{R}+z)\sigma_{-}$, (b) $H=(t_{L}+1/3z^{-2})\sigma_{+}+(t_{R}+z)\sigma_{-}$ (middle) and (c) $H=t_{L}\sigma_{+}+(0.6+0.12i+z+0.1z^{-2})\sigma_{-}$ (right), all with weak inter-chain coupling $\delta=1.6\times 10^{-3}$ and $t_{L},t_{R}=1.2\text{e}^{\pm 0.3}$ (here, $L$ increases from top to bottom). (d) While their GBZs can be irregularly-shaped, entanglement dips in $S$ consistently emerge at values of $L$ where the GBZ loop intersects the critical point $K_{0}$ (red).
Generality of entanglement dips.– SIEC and entanglement dips are not limited to the coupled $H_{\text{SSH}}$ chains explicitly computed so far. They are expected to show up whenever the spectrum depends on $L$, and becomes gapless at an EP at a special $L\approx L_{c}$. This requires a scale-dependent GBZ that can be generically designed by coupling subsystems with competing NHSE pumping. Presented in Fig. 4 are other models containing oppositely-directed NHSE channels, albeit with lower symmetry. Evidently, they all exhibit scale-dependent GBZs, and importantly exhibit entanglement dips at $L\approx L_{c}$ whenever the GBZ encounters a critical point $K_{0}$ (red).
@@
Physically, such dramatically suppressed EE represent the unique non-conservation of probability from antagonistic non-Hermitian pumping, which never occurs in ordinary NHSE processes where the gain/loss can be "gauged away" with a basis redefinition. Beyond the EE, entanglement dips also translate to kinks in the Renyi entropy, whose measurement prospects we discuss in I.
Acknowledgements.– J.G. acknowledges support by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant. C.H.L acknowledges support from the National Research Foundation, Singapore under its QEP2.0 programme (NRF2021-QEP2-02-P09), and the Ministry of Education, Singapore (MOE award number: MOE-T2EP50222-0003).
@@ I How to obtain the system size-dependent effective GBZ and spectrum
Recent work has revealed the crucial importance of the generalized Brillouin zones (GBZ) in describing non-Hermitian lattice Hamiltonians open boundary conditions (OBCs) [73, 150, 71]. For a given Hamiltonian $H(z)$ where $z=\exp(ik)$, the two roots $z$ of the characteristic (dispersion) equation $\text{det}(H(z)-E)=0$ with equal absolute values $|z|$ represents two “standing wave” solutions that can be superposed to satisfy OBCs at both ends. This is the typical way the GBZ is defined [96, 151, 69, 152, 153, 73, 99, 101, 154, 155, 156].
@@ I.1 Coupled 1D chains with opposite NHSE
We write down a model in which non-Hermitian 1D chains with antagonistic NHSE, denoted by $H$ and $H^{\dagger}$, are weakly coupled together by a small coupling parameter $\delta$. In the simplest cases, the coupling acts only between corresponding pairs of sites in a ladder-like fashion, such that the coupled Hamiltonian takes the form
$\begin{split}&\mathcal{H}=\begin{pmatrix}H&\delta\mathbb{I}\\ \delta\mathbb{I}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H^{T}}\end{pmatrix}\ ,\\ \end{split}$ (S1)
where $\delta\mathbb{I}$ is the identity matrix acting in the basis of $H$ or $H^{T}$, as shown in Fig. S1(a). We specialize $H$ to the well-known non-Hermitian SSH model [96, 97], which is simple enough for analytic solutions and yet topologically nontrivial:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E}=\pm\sqrt{(t_{L}+1/z)(t_{R}+z)}\ .$ (S3)
@@
Here, we first show how to compute the GBZ of fully uncoupled $H$ chains. We note that to satisfy OBCs at both ends of a 1D chain, an eigenfunction$|\Psi\rangle$ must be a linear combination [96] of at least two different non-Bloch wave functions $\sum_{n}z_{1}^{n}(|n,A\rangle+\phi_{z_{1}}|n,B\rangle)$ and $\sum_{n}z_{2}^{n}(|n,A\rangle+\phi_{z_{2}}|n,B\rangle)$, where $A,B$ and $n$ label the sublattices and unit-cells, $z_{1},z_{2}$ are the two solutions of Eq. S3 which satisfy $z_{1}z_{2}=t_{R}/t_{L}$, and $\phi_{z_{1,2}}=(t_{R}+z_{1,2})/E$. Crucially, the OBCs $\langle 0,B|\Psi\rangle=\langle L+1,A|\Psi\rangle=0$ enforce the constraint $z_{1}^{L+1}\phi_{z_{2}}=z_{2}^{L+1}\phi_{z_{1}}$ such that $(\phi_{z_{1}}/\phi_{z_{2}})^{1/L}\approx 1$. Consequently, for our uncoupled chain $H$, the GBZ satisfies $|z_{1}|=|z_{2}|$ and is simply given by $\left\{z=\sqrt{{t_{R}}/{t_{L}}}\exp(ik)|\,k=\frac{{2}m\pi}{L+1},m=1,2,\cdots,L\right\}$ i.e. with skin depth $\kappa_{H}=\frac{1}{2}\log|z|=\frac{1}{2}\log(t_{R}/t_{L})$. Its transpose conjugate is an oppositely directed chain $H^{T}$ that possesses skin depth $\kappa_{H^{\dagger}}=-\kappa_{H}=-\frac{1}{2}\log(t_{R}/t_{L})$.
@@ I.2 The new approach: Scale-dependent energies and GBZ
Instead, in our approach, we assign the decay parameter $z$ to one chain and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1/z}$ to the other chain. Ultimately, the OBC eigenstate would in general consist of a linear combination of two or more eigensolutions, which we will first solve for, separately. Writing each $z$ eigensolution as |$\ket{\varphi}=\sum_{n,\alpha}\varphi_{\alpha,n}\ket{n,\alpha}$ with sublattice $\alpha=A,B$(chain I), $A^{\prime},B^{\prime}$(chain II), we have $\varphi_{A(B),n}=\varphi_{B^{\prime}(A^{\prime}),L+1-n}$. Concretely,
@@
Due to spatial inversion symmetry ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}$, the decay in both chains are exactly equal but opposite – in more generic cases where the two coupled chains are not related by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}$, they would need to be separately solved for.
@@
$\left\{\begin{array}[]{ll}t_{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n-1}+\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A^{\prime},n}&=E{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n}\\ t_{R}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n+1}+\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B^{\prime},n}&=E{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n}\\ t_{R}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B^{\prime},n}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B^{\prime},n-1}+\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n}&=E{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A^{\prime},n}\\ t_{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A^{\prime},n}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A^{\prime},n+1}+\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n}&=E{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B^{\prime},n}\end{array}\right.\quad\stackrel{{\scriptstyle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}}}{{\longrightarrow}}\left\{\begin{array}[]{ll}t_{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n-1}+\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,L+1-n}&=E{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n}\\ t_{R}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,n+1}+\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{A,L+1-n}&=E{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi}_{B,n}\\ \end{array}\right.,$ (S12)
with site $n=2,3,\cdots,L-1$ and eigenvalue $E$. Due to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}$ symmetry, the four equations (from two chains with two sublattices each) are simplified to two equations. Substituting Eq. S5 into the above,
@@
For step (i), we work in the $\delta\ll 1$ limit and neglect the inter-chain couplings $\delta$ to obtain
@@
Via the same steps as in the previous subsection on a single chain, these bulk equations give (Eq. S3)
@@
|$\ket{\Psi_{\text{GBZ}}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|\varphi_{z_{1}}\rangle-c|\varphi_{z_{2}}\rangle}=\sum^{L}_{n=1}\sum_{\alpha=A,B,A^{\prime},B^{\prime}}\psi_{n,\alpha}|n,\alpha\rangle=\sum_{n}\left(\begin{pmatrix}z_{1}^{n}\\ z_{1}^{n}\phi_{z_{1}}\\ z_{1}^{L+1-n}\phi_{z_{1}}\\ z_{1}^{L+1-n}\end{pmatrix}-c\begin{pmatrix}z_{2}^{n}\\ z_{2}^{n}\phi_{z_{2}}\\ z_{2}^{L+1-n}\phi_{z_{2}}\\ z_{2}^{L+1-n}\end{pmatrix}\right)\ket{n},$ (S23)
where the coefficient $c$ controls the relative amplitude of the $z_{1},z_{2}$ non-Bloch contributions. We next substitute this ansatz Eq. S23 into the open boundary conditions, where $\psi_{n,\alpha}$ disappear at $n=0$ and $L+1$ (and beyond):
$\left\{\begin{array}[]{ll}t_{L}\psi_{B,1}+\delta\psi_{A^{\prime},1}&=E\psi_{A,1}\\ t_{R}\psi_{A,L}+\delta\psi_{B^{\prime},L}&=E\psi_{B,L}\\ t_{R}\psi_{B^{\prime},1}+\delta\psi_{A,1}&=E\psi_{A^{\prime},1}\\ t_{L}\psi_{A^{\prime},L}+\delta\psi_{B,L}&=E\psi_{B^{\prime},L}\end{array}\right.\stackrel{{\scriptstyle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{I}}}}{{\longrightarrow}}\left\{\begin{array}[]{ll}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta\psi_{B,L}=E\psi_{A,1}-t_{L}\psi_{B,1}}&\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta\psi_{A,1}=E\psi_{B,L}-t_{R}\psi_{A,L}}&\\ \end{array}\right..$ (S30)
To proceed, the key idea is that the extra terms involving the small inter-chain coupling $\delta$, which were not taken into account in deriving the bulk $z_{1},z_{2}$ solutions, must exactly compensate the missing terms due to the open boundaries. By explicitly substituting $E$ in terms of $z_{1,2}$, $\phi_{z_{1,2}}$ and the hoppings using Eq. S19, we obtain
@@
We comment that $z_{1},z_{2}$, which were previously derived under the uncoupled ($\delta=0$) approximation, has now become dependent on the coupling $\delta$ through $E$ (even though this dependence is hidden behind Eq. S19, which was derived under the uncoupled approximation). Since the inter-chain couplings are weak, we are safe to assume that they cannot fundamentally reverse the NHSE direction within each chain. That is, in the case of $|t_{R}/t_{L}|<1$, we have $|z_{2}|\leq\sqrt{|t_{R}/t_{L}|}\leq|z_{1}|\leq 1$. Together with $\delta\ll 1$ and $|z_{j}|<1$, we can thus approximate $1-\delta z_{j}^{L}\approx 1$ ($j=1,2$), thereby simplifying the above to
@@
Below, we highlight two extreme but highly relevant cases:
  • $\delta\ll{t_{R}^{L}}/{t_{L}^{L}}$, which is the regime where the coupling $\delta$ is so weak that it remains negligible despite the presence of NHSE competition. In this case, only the second term under the square root in Eq. S38 survives:
@@
    As $L+1]{{\phi_{z_{1}}}/{\phi_{z_{2}}}}\approx 1$ for moderately large $L$, where $\phi_{z}=(t_{R}+z)/E$, we recover
@@
  • $\delta\gg{t_{R}^{L}}/{t_{L}^{L}}$, which defines the strongly coupled regime. Here, the first term under the square root in Eq. S38 dominates, such that

### v2

@@ Abstract
which, unlike in conventional critical scenarios of Figs. 1b,c, diverges due to $L\rightarrow L_{c}$ rather than vanishing momentum spacings. This square-root divergence shows up in the $\bar{P}$ spectrum near $L\approx L_{c}$ [Fig. 3a, also see the Supplementary material] which in turn translates to a characteristic $S\sim 1/\sqrt{|L_{c}-L|}$ divergence in the EE which we call the entanglement dip [Fig. 3b].
While such entanglement dips can in principle be infinitely deep, in practice its depth is limited by how closely $L_{c}$ approaches an integer. Shown in Fig. 3c is the minimal entanglement $S_{\text{min}}$ as the hopping asymmetry $t_{L}/t_{R}$ and inter-chain coupling $\delta$ are varied. For each lattice size $L$, the EE dip becomes particularly deep along curves in parameter space (dark pink), even reaching $S_{\text{min}}\approx-3$ in the region shown.
Figure 4: Entanglement dips and scale-dependent GBZ in more general models satisfying Eq. (7). Shown are the $L$-dependent GBZs for illustrative (a) $H=(t_{L}+1/z)\sigma_{+}+(t_{R}+z)\sigma_{-}$, (b) $H=(t_{L}+1/3z^{-2})\sigma_{+}+(t_{R}+z)\sigma_{-}$ (middle) and (c) $H=t_{L}\sigma_{+}+(0.6+0.12i+z+0.1z^{-2})\sigma_{-}$ (right), all with weak inter-chain coupling $\delta=1.6\times 10^{-3}$ and $t_{L},t_{R}=1.2\text{e}^{\pm 0.3}$ (here, $L$ increases from top to bottom). (d) While their GBZs can be irregularly-shaped, entanglement dips in $S$ consistently emerge at values of $L$ where the GBZ loop intersects the critical point $K_{0}$ (red).
Generality of entanglement dips.– SIEC and entanglement dips are not limited to the coupled $H_{\text{SSH}}$ chains explicitly computed so far. They are expected to show up whenever the spectrum depends on $L$, and becomes gapless at an EP at a special $L\approx L_{c}$. This requires a scale-dependent GBZ that can be generically designed by coupling subsystems with competing NHSE pumping. Presented in Fig. S13 are other models containing oppositely-directed NHSE channels, albeit with lower symmetry. Evidently, they all exhibit scale-dependent GBZs, and importantly exhibit entanglement dips at $L\approx L_{c}$ whenever the GBZ encounters a critical point $K_{0}$ (red).
@@
Physically, such dramatically suppressed EE represent the unique non-conservation of probability from antagonistic non-Hermitian pumping, which never occurs in ordinary NHSE processes where the gain/loss can be "gauged away" with a basis redefinition. Beyond the EE, entanglement dips also translate to kinks in the Renyi entropy, whose measurement prospects we discuss in Supplementary material Section V.
Conflict of interest. The authors declare that they have no conflict of interest.
Acknowledgements. J.G. acknowledges support by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant. C.H.L acknowledges support from the Ministry of Education, Singapore (MOE award number: MOE-T2EP50222-0003). H.J is supported by “the Fundamental Research Funds for the Central Universities.
Author contributions. Ching Hua Lee proposed the initial idea and led the whole project. Hui Jiang co-led the project and designed key calculations. Sirui Liu and Hui Jiang executed the project and conducted key calculations. Wen-Tan Xue and Qingya Li performed numerical simulations and data analysis. Jiangbin Gong and Xiaogang Liu participated in the discussions and supervised the research. All participated in the writing of the manuscript.
@@ I How to obtain the system size-dependent effective GBZ and spectrum
Recent work has revealed the crucial importance of the generalized Brillouin zones (GBZ) in describing non-Hermitian lattice Hamiltonians open boundary conditions (OBCs) [1, 2, 3, 4, 5]. For a given Hamiltonian $H(z)$ where $z=\exp(ik)$, the two roots $z$ of the characteristic (dispersion) equation $\text{det}(H(z)-E)=0$ with equal absolute values $|z|$ represents two “standing wave” solutions that can be superposed to satisfy OBCs at both ends. This is the typical way the GBZ is defined [6, 7, 8, 9, 10, 1, 11, 12, 13, 14, 15].
@@ I.1 Coupled 1D chains with opposite NHSE
We write down a model in which non-Hermitian 1D chains with antagonistic NHSE, denoted by $H$ and $H^{T}$, are weakly coupled together by a small coupling parameter $\delta$. In the simplest cases, the coupling acts only between corresponding pairs of sites in a ladder-like fashion, such that the coupled Hamiltonian takes the form
$\begin{split}&\mathcal{H}=\begin{pmatrix}H&\delta\mathbb{I}\\ \delta\mathbb{I}&H^{T}\end{pmatrix}\ ,\\ \end{split}$ (S1)
where $\delta\mathbb{I}$ is the identity matrix acting in the basis of $H$ or $H^{T}$, as shown in Fig. S1a. We specialize $H$ to the well-known non-Hermitian SSH model [6, 16], which is simple enough for analytic solutions and yet topologically nontrivial:
@@
${E}=\pm\sqrt{(t_{L}+1/z)(t_{R}+z)}\ .$ (S3)
@@
Here, we first show how to compute the GBZ of fully uncoupled $H$ chains. We note that to satisfy OBCs at both ends of a 1D chain, an eigenfunction$|\Psi\rangle$ must be a linear combination [6] of at least two different non-Bloch wave functions $\sum_{n}z_{1}^{n}(|n,A\rangle+\phi_{z_{1}}|n,B\rangle)$ and $\sum_{n}z_{2}^{n}(|n,A\rangle+\phi_{z_{2}}|n,B\rangle)$, where $A,B$ and $n$ label the sublattices and unit-cells, $z_{1},z_{2}$ are the two solutions of Eq. (S3) which satisfy $z_{1}z_{2}=t_{R}/t_{L}$, and $\phi_{z_{1,2}}=(t_{R}+z_{1,2})/E$. Crucially, the OBCs $\langle 0,B|\Psi\rangle=\langle L+1,A|\Psi\rangle=0$ enforce the constraint $z_{1}^{L+1}\phi_{z_{2}}=z_{2}^{L+1}\phi_{z_{1}}$ such that $(\phi_{z_{1}}/\phi_{z_{2}})^{1/L}\approx 1$. Consequently, for our uncoupled chain $H$, the GBZ satisfies $|z_{1}|=|z_{2}|$ and is simply given by $\left\{z=\sqrt{{t_{R}}/{t_{L}}}\exp(ik)|\,k=\frac{{2}m\pi}{L+1},m=1,2,\cdots,L\right\}$ i.e. with skin depth $\kappa_{H}=\frac{1}{2}\log|z|=\frac{1}{2}\log(t_{R}/t_{L})$. Its transpose conjugate is an oppositely directed chain $H^{T}$ that possesses skin depth $\kappa_{H^{\dagger}}=-\kappa_{H}=-\frac{1}{2}\log(t_{R}/t_{L})$.
@@ I.2 The new approach: Scale-dependent energies and GBZ
Instead, in our approach, we assign the decay parameter $z$ to one chain and ${1/z}$ to the other chain. Ultimately, the OBC eigenstate would in general consist of a linear combination of two or more eigensolutions, which we will first solve for, separately. Writing each $z$ eigensolution as |$\ket{\varphi}=\sum_{n,\alpha}\varphi_{\alpha,n}\ket{n,\alpha}$ with sublattice $\alpha=A,B$(chain I), $A^{\prime},B^{\prime}$(chain II), spatial inversion symmetry $\mathcal{I}$ enforces $\varphi_{A(B),n}=\varphi_{B^{\prime}(A^{\prime}),L+1-n}$. Concretely,
@@
Due to spatial inversion symmetry ${\mathcal{I}}$, the decay in both chains are exactly equal but opposite – in more generic cases where the 2 coupled chains are not related by spatial inversion symmetry ${\mathcal{I}}$, they would need to be separately solved for.
@@
$\left\{\begin{array}[]{ll}t_{L}{\varphi}_{B,n}+{\varphi}_{B,n-1}+\delta{\varphi}_{A^{\prime},n}&=E{\varphi}_{A,n}\\ t_{R}{\varphi}_{A,n}+{\varphi}_{A,n+1}+\delta{\varphi}_{B^{\prime},n}&=E{\varphi}_{B,n}\\ t_{R}{\varphi}_{B^{\prime},n}+{\varphi}_{B^{\prime},n-1}+\delta{\varphi}_{A,n}&=E{\varphi}_{A^{\prime},n}\\ t_{L}{\varphi}_{A^{\prime},n}+{\varphi}_{A^{\prime},n+1}+\delta{\varphi}_{B,n}&=E{\varphi}_{B^{\prime},n}\end{array}\right.\quad\stackrel{{\scriptstyle{\mathcal{I}}}}{{\longrightarrow}}\left\{\begin{array}[]{ll}t_{L}{\varphi}_{B,n}+{\varphi}_{B,n-1}+\delta{\varphi}_{B,L+1-n}&=E{\varphi}_{A,n}\\ t_{R}{\varphi}_{A,n}+{\varphi}_{A,n+1}+\delta{\varphi}_{A,L+1-n}&=E{\varphi}_{B,n}\\ \end{array}\right.,$ (S12)
with site $n=2,3,\cdots,L-1$ and eigenvalue $E$. Due to ${\mathcal{I}}$ symmetry, the four equations (from two chains with two sublattices each) are simplified to two equations. Substituting Eq. (S5) into the above,
@@
For step (i), we work in the $\delta\ll 1$ limit and neglect the inter-chain couplings $\delta$. In practice, it is sufficient for $\delta$ to be at least one order of magnitude smaller than other system parameters. This yields the decoupled eigenequation to obtain
@@
Via the same steps as in the previous subsection on a single chain, these bulk equations give (Eq. (S3))
@@
|$\ket{\Psi_{\text{GBZ}}}={|\varphi_{z_{1}}\rangle-c|\varphi_{z_{2}}\rangle}=\sum^{L}_{n=1}\sum_{\alpha=A,B,A^{\prime},B^{\prime}}\psi_{n,\alpha}|n,\alpha\rangle=\sum_{n}\left(\begin{pmatrix}z_{1}^{n}\\ z_{1}^{n}\phi_{z_{1}}\\ z_{1}^{L+1-n}\phi_{z_{1}}\\ z_{1}^{L+1-n}\end{pmatrix}-c\begin{pmatrix}z_{2}^{n}\\ z_{2}^{n}\phi_{z_{2}}\\ z_{2}^{L+1-n}\phi_{z_{2}}\\ z_{2}^{L+1-n}\end{pmatrix}\right)\ket{n},$ (S23)
where the coefficient $c$ controls the relative amplitude of the $z_{1},z_{2}$ non-Bloch contributions. We next substitute this ansatz Eq. (S23) into the open boundary conditions, where $\psi_{n,\alpha}$ disappear at $n=0$ and $L+1$ (and beyond):
$\left\{\begin{array}[]{ll}t_{L}\psi_{B,1}+\delta\psi_{A^{\prime},1}&=E\psi_{A,1}\\ t_{R}\psi_{A,L}+\delta\psi_{B^{\prime},L}&=E\psi_{B,L}\\ t_{R}\psi_{B^{\prime},1}+\delta\psi_{A,1}&=E\psi_{A^{\prime},1}\\ t_{L}\psi_{A^{\prime},L}+\delta\psi_{B,L}&=E\psi_{B^{\prime},L}\end{array}\right.\stackrel{{\scriptstyle{\mathcal{I}}}}{{\longrightarrow}}\left\{\begin{array}[]{ll}{\delta\psi_{B,L}=E\psi_{A,1}-t_{L}\psi_{B,1}}&\\ {\delta\psi_{A,1}=E\psi_{B,L}-t_{R}\psi_{A,L}}&\\ \end{array}\right..$ (S30)
To proceed, the key idea is that the extra terms involving the small inter-chain coupling $\delta$, which were not taken into account in deriving the bulk $z_{1},z_{2}$ solutions, must exactly compensate the missing terms due to the open boundaries. By explicitly substituting $E$ in terms of $z_{1,2}$, $\phi_{z_{1,2}}$ and the hoppings using Eq. (S19), we obtain
@@
We comment that $z_{1},z_{2}$, which were previously derived under the uncoupled ($\delta=0$) approximation, has now become dependent on the coupling $\delta$ through $E$ (even though this dependence is hidden behind Eq. (S19), which was derived under the uncoupled approximation). Since the inter-chain couplings are weak and the spatial inversion symmetry, we are safe to assume that they cannot fundamentally reverse the NHSE direction within each chain. That is, in the case of $|t_{R}/t_{L}|<1$, we have $|z_{2}|\leq\sqrt{|t_{R}/t_{L}|}\leq|z_{1}|\leq 1$. Together with $\delta\ll 1$ and $|z_{j}|<1$($j=1,2$), note that for small systems $|z_{j}|$ is well below 1, whereas for large systems $|z_{j}|$ may approach 1 yet $\lim_{L\gg 1}|z_{j}|^{L}$ still decays rapidly to zero. Moreover, the imaginary part of $z_{j}^{L}$ can be neglected since the effective model requires the momentum to satisfy periodic boundary conditions. Therefore, we can approximate $1-\delta z_{j}^{L}\approx 1$ ($j=1,2$), thereby simplifying the above to
@@
Below, we highlight 2 extreme but highly relevant cases:
  • $\delta\ll{t_{R}^{L}}/{t_{L}^{L}}$, which is the regime where the coupling $\delta$ is so weak and imposes a constraint on the system size, that it remains negligible despite the presence of NHSE competition. In this case, only the second term under the square root in Eq. (S38) survives
@@
    As $L+1]{{\phi_{z_{1}}}/{\phi_{z_{2}}}}\approx 1$ for moderately large $L$ ($L\geq 10$), where $\phi_{z}=(t_{R}+z)/E$, we recover
@@
  • $\delta\gg{t_{R}^{L}}/{t_{L}^{L}}$, which defines the strongly coupled regime. Here, the first term under the square root in Eq. (S38) dominates, such that